If 5000 dollars is invested in a bank account at an interest rate of 4 per cent per year, compounded continuously. How many years will it take for your balance to reach 30000 dollars?
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

If 5000 dollars is invested in a bank account at an interest rate of 4 per cent per year, compounded continuously. How many years will it take for your balance to reach 30000 dollars?
----
A(t) = Pe^(rt)
30000 = 5000*e^(0.04t)
6 = 3^(0.04t)
---
Use natural logs to find "t":
0.04t*ln(3) = ln(6)
---
0.04t = ln(6)/ln(3)
---
0.04t = 1.0631
---
t = 40.8 years
====================
Cheers,
Stan H.
====================

Answer by solver91311(24713)



A very long time.



Substitute values





Take the natural log of both sides



Use the laws of logs and the fact that



years





John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism